OpenStudy (anonymous):

Write an equation of a line parallel to line GH below in slope-intercept form that passes through the point (-5, 6). Line GH is shown. G is at negative 2, 2. H is at 2, 6. y = −x − 1 y = x − 1 y = −x + 11 y = x + 11

OpenStudy (anonymous):

A. y = -x - 1 B. y = x - 1 C. y = -x + 11 D. y = x + 11

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@tanya123

OpenStudy (anonymous):

@jim_thompson5910